# SR-CACO-2: A Dataset for Confocal Fluorescence Microscopy Image Super-Resolution

### v1

@@ 3.1 Dataset Capture:
3.2 Experimental Protocol:
@@
Patches sampling. Since the imaged tiles are quite large ( ${\leavevmode\nobreak\ 9k\times 9k}$), they are not well suited for machine learning models as they usually operate on smaller images. To this end, we pre-processed the tiles to patches as presented in Fig.4. In particular, on HR tiles of the brightest cell, i.e. CELL2, we used a sliding window that scans from left to right and from top to bottom of the entire tile. Windows can overlap by ${25\%}$ to increase the chance of capturing large patterns and minimize cropping cells. The window size is set to ${576\times 576}$ with 6464{64}64 extra margin at each side that will be useful for a later operation (registration). This additional margin was discarded, and only windows of size ${512\times 512}$ were preserved as final patches.
@@
Many of the sampled windows will land on empty regions, i.e., black background. In particular, across all the 22 HR tiles of CELL2, ${74.24\%}$ of total pixels are background while only ${25.74\%}$ are foreground (cells). Windows which are more black will not be helpful in learning or evaluation since they are easy to reconstruct using models. Therefore, they will introduce an evaluation bias by giving a high level of performance, which does not reflect the true performance of the models to super-resolve relevant regions, i.e., cells. For this reason, we discarded the windows without enough cell content. To detect these windows, we threshold the tile using a threshold value of 44{4}4 which is manually deemed sufficient to spot the cells without triggering tiny noise to appear. Only windows with at least ${20\%}$ of cell content were preserved.
@@
Data split. First, the 22 tiles are randomly split into 15 for training, 3 for validation, and 4 for testing. This prevents mixing patches of the same tile across different sets. Splitting the tiles is done randomly. Tiles with identifiers (ID) = [’9’, ’10’, ’14’, ’20’] are assigned to the test set. Tiles with ID [’7’, ’11’, ’19’] are used for the validation set. Finally, tiles with ID = [’1’, ’2’, ’3’, ’4’, ’5’, ’6’, ’8’, ’12’, ’13’, ’15’, ’16’, ’17’, ’18’, ’21’, ’22’] are used for training. This is summarized in Tab.1 along with the patch distribution across all subsets.
@@ 4 Single-Image Super-Resolution Methods
In addition to the proposed new dataset, we provide in this work an initial benchmarking of different SISR methods. In particular, we evaluated 15 SISR methods taken from 4 common families [70]. In addition, we compare to a simple baseline which is interpolation via bicubic method. The reported results here can be used as baselines for future comparisons. We select representative and most recent methods as follows: Pre-upsampling SR: SRCNN [14], VDSR [30], DRRN [58], and MemNet [59]. Post-upsampling SR: NLSN [42], DFCAN [47], SwinIR [40], ENLCN [73], GRL [38], ACT [74], and Omni-SR [65]. Iterative up-and-down sampling SR: DBPN [20], and SRFBN [39]. Progressive upsampling SR: ProSR [67], and MS-LapSRN [32].
We note that post-upsampling family is the most dominant approach [70]. It contains most recent works. Other families have seen less progress mainly due to their high computational cost.
@@ 4.1 Evaluation Metrics
The evaluation is performed on the full patch, referred to as full image, as commonly done. We furthermore report refined performance over the cells only, referred to as ROI only, to better assess the prediction quality without the black background inside the patches. ROIs are determined by thresholding the HR patch using a set of thresholds [^2] Evaluation thresholds are [4, 5, 6, 7, 8, 9, 10].. A performance is computed per-threshold over the ROI only. The final reported performance is the average of all per-threshold performances. Models are trained on each cell type, and each scale separately.
@@ 4.2 Implementation Details
SISR task: Tables 4, 5 show results over all the methods on ROI only, i.e., cells, and on full image, respectively. Figure 6 shows the PSNR performance in a visual form across all the cells, scales, and methods. Note that performance over full images are relatively way higher compared to when only ROI is considered. Overall, SISR methods achieved better results than simple interpolation. As commonly known, the higher the scale factor, the more difficult the task. This is also observed in our results across all methods and performance measures. For instance, SRCNN [14] yields a PSNR of 35.0835.08{35.08}35.08, 33.1733.17{33.17}33.17, and 29.1929.19{29.19}29.19 for X2, X4, and X8, respectively. Another emerging pattern in these results is the discrepancy performance across cell types: CELL0, CELL1, and CELL2. Over PSNR and SSIM performance, CELL0, i.e. dimmest, is the easiest while CELL2 is the most difficult, i.e. the brightest. Over NRMSE, results are mixed. Across all the four studied SISR families, we observe that pre-upsampling SR yielded better results overall across all metrics and scales. This is unexpected since post-upsampling SR are current state-of-the-art approaches [70]. These results may suggest that upscaling our LR images through deep models may fail to reconstruct the details of our HR images. This may be explained by the nature of the details present in our HR images which paradoxically comes from noise. Confocal microscopes perform multiple scans to produce a set of slices that are aggregated to produce a final 2D image. Due to an inherited noise produced by the detectors on microscopes, each slice is different. However, averaging these slices decreases the noise and maintains the true signal. The work of [63] shows that it is difficult for CNN based models to reconstruct input images with noise as they tend to filter out that noise even when it is provided as input. Therefore, it is challenging for such models to produce similar images to our HR images. Such result should be considered in designing future SISR methods for this dataset. We provide in Fig.7 typical visual results for scale X2 over CELL2 which illustrate the limited capacity of these methods to produce accurate HR details as they yield blurry images.
@@
Downstream biology tasks: Object detection/segmentation. To evaluate the image results from each model on downstream biology tasks, we evaluated each model at X2, X4, and X8 using an object segmentation task using widely used and openly accessible software for image analysis (ImageJ) [52] [^3] https://imagej.net/ij. 30 random images from test set are selected. Images were resized to 512 x 512 pixels with no interpolation. Automatic object segmentation was performed using Star-convex Shapes (StarDist) [51, 71, 72] with the following parameters: modelChoice = Versatile (fluorescent nuclei), normalizeInput = TRUE, percentileBottom = 1.1, percentileTop = 100.0, probThresh = 0.6, nmsThresh = 0.4. The resulting output was the number of objects detected. Manual object counting was performed to create the Ground Truth number of objects. Manual comparisons were made to identify oversegmentation and under segmentation errors for each image. Graphing and statistical analysis were performed using JMP Pro (version 17.0.0) [^4] https://www.jmp.com/en_ca/home.html. Steel’s test for non-parametric multiple comparisons using HR as the control group was performed to assess model performance. Our objective was to evaluate model performance relative to HR, not ground truth. Models that perform as well as HR (i.e. not statistically different than HR at p=0.05) are represented in orange. Figures 8, 9 show results for cell detection across the 2 extreme cases: CELL2 and CELL0, respectively. Cell segmentation performance is reported in Fig.10. Overall, several methods (orange) achieved comparable results to HR images on these tasks which is promising. Figure 11 shows an example of cell segmentation of different methods.
5 Ethical Considerations and Dataset Accessibility
The SR-CACO-2 dataset does not require ethics approval. Caco-2 (Cancer coli-2) cells were isolated from a 72 year old white male in the 1970s as part of a collection of gastrointestinal cancer cell lines at the Memorial Sloan Kettering Cancer Center [34]. The use of Caco-2 cells does not require ethics approval because the cells were established prior to the US Federal Policy for the Protection of Human Subjects (1991) and the UK Human Tissue Act (2004). This is the standard for the biological field for the use of cell lines from derived from human tissue that are now considered publicly available. The cells were obtained from the American Type Culture Collection (ATCC) [^5] https://www.atcc.org, a nonprofit organization that collects and distributes cell lines and other biological materials that are publicly accessible. Its legal disclaimers allow the use of Caco-2 cells for laboratory research https://www.atcc.org/products/htb-37.
The dataset is made public and can be freely accessed for academic, non-commercial and not-for-profit research under a custom license. A user is required sign an end-user license agreement to request the dataset. Once approved, we send them the link to download the dataset. The dataset which includes full tiles, and patches, are stored permanently on Google Drive. The code, along with the pretrained weights use an open-source license and they are located at https://github.com/sbelharbi/sr-caco-2, and https://huggingface.co/sbelharbi/sr-caco-2, respectively.
6 Dataset Publishing and Usage
6.1 Dataset Hosting and Licensing
The dataset and code are available for researchers. We use a custom non-commercial license for research purposes only. To access the dataset, a user is required to sign an end-user license agreement. Upon approval, a link to the dataset is communicated to the user. The dataset represents 1.8 GB for all the tiles, and 3.4 GB for all the patches. The dataset is stored in Google Drive  [^6] https://drive.google.com intended for long time availability. The Google Drive account is accessible by the all the authors of the dataset for future updates, and withdrawal if needed. The Github code web-page is meant as the front-end site for the dataset https://github.com/sbelharbi/sr-caco-2. The ReamMe file contains a brief overview of the dataset, request instructions, link to a permanent arXiv [^7] https://arxiv.org pdf with the content of this paper, link to download the pretrained weights, and installation and usage of the code. The code uses an open-source 3-Clause BSD License [1]. Due to their large size, the pretrained weights of different benchmarked methods are stored in Hugging Face https://huggingface.co/sbelharbi/sr-caco-2 under the same license as the code.
6.2 Intended Uses and Ethical Considerations
The dataset is made available for research purposes only. No personal information of human subject who provided the cells is available. Based on our dataset, it is impossible to uncover their identity.
The dataset is meant primarily to train models for SISR task for confocal fluorescence microscopy imaging for 3 different scales (X2, X4, and X8). Accurate models trained with our dataset can be efficiently used to performance SISR on live-imaging of low-resolution videos. This allows fast imaging at low-resolution, and therefore, reduce the cell damage caused by long exposure to light, and also allowing the observation of instantaneous inter-cellular events. Additionally, improving live-imaging videos quality has a huge benefit for downstream biological tasks including cell segmentation, cell counting, cell movement tracking, identifying and characterizing cell divisions.
Provided the adequate annotation, this dataset could be used for other biological/cell related applications for computational training such as: Cell segmentation: the availability of both nuclei and cell membranes makes cell segmentation possible. However, the dataset does not exhibit multiple different cell types/cell shapes, making it less universal. Cell count: Models can be trained to identify cell number, cell shape, perhaps also identifying dividing cells as well, since there is staining for chromosomes as well as Tubulin/Survivin, which both mark certain stages of division.
7 SR-CACO-2 Dataset Limitations
First, it would have been interesting to supplement this data set with oversaturation/undersaturation or very bright vs very dim imaging, as live imaging of cells requires low laser power to minimize damage to cells. This therefore often results in dim images, compared to what can be acquired using fixed and immunostained images. Second, Caco2 cells were ideal for cell for use in this study due to the ease and efficiency of culturing them in 3D and their ability to efficiently polarize. However, often the next step in biological studies is to use human patient-derived tissue or mouse tissue. Future studies could involve the imaging of either primary human or primary mouse organoids. Third, live cells are more dynamic, while fixed cells are static. Live cells are actively growing, dividing, and synthesizing building blocks. This results in changes to the markers that we have used, like mCherry-Histone H2B and GFP-tubulin, which can change shape and position within an individual cell over time, and provide important biological information, such as the stage of cell division. This information cannot be followed for an individual cell once fixed and immunostained, which may impact the SISR model performance. Although, live cells look nearly identical to fixed cells, using fixed dead cells for training SISR models remains a practical solution. Ideally, collecting images of live cells as time-lapse videos for training could be better choice as it reflects the test time scenario. However, acquiring such data is extremely challenging for the task of SISR due to cell movements or appearance of new cells due to cell division. This will create a large misalignment between tiles at different scales, rendering the training impractical. Forth, the dataset is limited in term of the types of proteins that have been captured. We are only looking at chromosomes, cell membrane/cell structural proteins, and proteins specific for division. The dataset could perhaps be improved with an increase in the number and type of markers used, such as proteins that appear as puncta in the cytoplasm, proteins localized to the apical membrane (mutually exclusive localization with E-cadherin), basement membrane proteins (extracellular matrix found around the outside of cysts), and signaling molecules.

### v2

@@ 3.1 Dataset Capture
3.2 Dataset Variability
During the acquisition of SR-CACO-2, we ensured the technical and biological diversity of samples. To automate the image acquisition, we programmed the microscope to capture multiple sets of 100 images. Each set of 100 images was stitched together (${10\times 10}$) to form an image tile. Therefore, each tile represents 100 unique images (fields of view) and the 22 tiles correspond to 2,200 unique images. Each image is captured at four scales and for three markers. The automatic collection of a tile of the 4 different resolutions takes over $\sim$12-16 hours. We note that each tile was captured in a multi-well plate. The 22 wells were collected from 4 independent experiments using 5-6 wells per plate/experiment, ensuring adequate and standard biological diversity.
Figure 5: Object-based analysis of cellular structures captured of all the 2,200 high-resolution images (${22\times 10\times 10}$).
To further assess the diversity of our dataset, we performed an object-based analysis of cellular structures captured in the image dataset. The 2,200 (${22\times 10\times 10}$) unique image fields of view at the high resolution contain ${\sim 16,800}$ multi-cellular objects. The cells were cultivated in a three-dimensional protein matrix that allows them to form tissue-like cyst structures that resemble the natural organization of cells in tissues and organs, such as the colon, lung, breast, etc. Cells grown in a three-dimensional matrix exhibit spontaneous variations in cyst size and organization (hollow or solid). We recently [87, 86] reported multiple cellular phenotypes in Caco2 cells that resemble tissue geometries found in healthy and disease states. Our dataset contains cysts with size variations over 2 orders of magnitude. Moreover, shape analysis of the cysts demonstrates substantial variation in aspect ratio and circularity. This reflects cysts covering a spectrum of phenotypes, including single-layered, multi-layered, and solid. This contrasts with cells grown as two-dimensional cultures on plastic or glass surfaces, where there is substantially less variation in size and organization when viewing single cells or confluent patches as is often used in other recently published microscopy datasets. Fig.5 shows the analysis of the cells in the SR-CACO-2 dataset. We can observe from these plots that the cells span a large size range with a relatively circular shape.
Our dataset also includes diverse cellular markers with unique properties. Cells were labeled with Hoechst dye that labels chromatin and is frequently used as a nuclear stain in cell biology. We also include mCherry-Histone, a red-fluorescent protein conjugated to histone H2B that is part of chromatin. This serves as an additional marker of cell nuclei and represents a signal that is often employed in live imaging experiments. mCherry-Histone represents a lower signal than Hoechst because the cells were selected to have a low-to-moderate expression of this protein and are detected by direct fluorescence. The other two markers (E-cadherin and Survivin) represent typical staining obtained with indirect fluorescence methods, using primary and fluorophore-conjugated secondary antibodies. E-cadherin is expressed in all Caco2 cells and is a cell surface protein. Therefore, this has a distinct expression pattern from the blob-like nuclear labels described above. Despite being expressed in all cells, there is heterogeneity at the cellular level, providing extensive natural variation within our dataset. Survivin is transiently expressed in cells during cell division. This represents an additional unique marker pattern that appears as an intense patch connecting two cells. Survivin is only expressed in a subset of cells during mitosis and cytokinesis, processes that often utilize imaging and machine learning for analysis [40].
3.3 Ethical Considerations and Dataset Accessibility
The SR-CACO-2 dataset does not require ethics approval. Caco-2 (Cancer coli-2) cells were isolated from a 72-year-old white male in the 1970s as part of a collection of gastrointestinal cancer cell lines at the Memorial Sloan Kettering Cancer Center [46]. The use of Caco-2 cells does not require ethics approval because the cells were established prior to the US Federal Policy for the Protection of Human Subjects (1991) and the UK Human Tissue Act (2004). This is the standard for the biological field for the use of cell lines derived from human tissue that are now considered publicly available. The cells were obtained from the American Type Culture Collection (ATCC) [^1] https://www.atcc.org, a nonprofit organization that collects and distributes cell lines and other biological materials that are publicly accessible. Its legal disclaimers allow for the use of Caco-2 cells for laboratory research https://www.atcc.org/products/htb-37.
The dataset is made public and can be freely accessed under the Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International license (CC BY-NC-SA 4.0) https://creativecommons.org/licenses/by-nc-sa/4.0/. The dataset which includes full tiles, and patches, is stored permanently on Google Drive. The provided code uses an open-source license and it is located at https://github.com/sbelharbi/sr-caco-2 along with the dataset downloading link. The pretrained weights can be found at https://huggingface.co/sbelharbi/sr-caco-2.
3.4 Dataset Hosting and Licensing
The dataset and code are available for researchers. The dataset is released under a Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International license (CC BY-NC-SA 4.0) https://creativecommons.org/licenses/by-nc-sa/4.0/. It is freely accessible. The dataset represents 1.8 GB for all the tiles, and 3.4 GB for all the patches. The dataset is stored in Google Drive  [^2] https://drive.google.com intended for long-time availability. The Google Drive account is accessible by the all authors of the dataset for future updates, and withdrawal if needed. The Github code web page is meant as the front-end site for the dataset https://github.com/sbelharbi/sr-caco-2. The ReamMe file contains a brief overview of the dataset, its download link, link to a permanent arXiv [^3] https://arxiv.org pdf with the content of this paper, link to download the pretrained weights, and installation and usage of the code. The code uses an open-source 3-Clause BSD License [1]. Due to their large size, the pretrained weights of different benchmarked methods are stored in Hugging Face https://huggingface.co/sbelharbi/sr-caco-2 under the same license as the code.
3.5 Intended Uses and Ethical Considerations
The SR-CACO-2 dataset does not require ethics approval. Caco-2 (Cancer coli-2) cells were isolated from a 72-year-old white male in the 1970s as part of a collection of gastrointestinal cancer cell lines at the Memorial Sloan Kettering Cancer Center [46]. The use of Caco-2 cells does not require ethics approval because the cells were established before the US Federal Policy for the Protection of Human Subjects (1991) and the UK Human Tissue Act (2004). This is the standard for the biological field to use cell lines from humans that are now considered publicly available. The cells were obtained from the American Type Culture Collection (ATCC) [^4] https://www.atcc.org, a nonprofit organization that collects and distributes cell lines and other biological materials that are publicly accessible.
The dataset is made available freely under the Creative Commons license CC BY-NC-SA 4.0. No personal information of human subjects who provided the cells is available. Based on our dataset, it is impossible to uncover their identity.
The dataset is meant primarily to train models for SISR tasks for confocal fluorescence microscopy imaging for 3 different scales (X2, X4, and X8). Accurate models trained with our dataset can be efficiently used to perform SISR on live imaging of low-resolution videos. This allows fast imaging at low resolution, and therefore, reduces the cell damage caused by long exposure to light, and also allows the observation of instantaneous inter-cellular events. Additionally, improving live-imaging video quality has a huge benefit for downstream biological tasks including cell segmentation, cell counting, cell movement tracking, and identifying and characterizing cell divisions.
Provided the adequate annotation, this dataset could be used for other biological/cell-related applications for computational training such as: Cell segmentation: the availability of both nuclei and cell membranes makes cell segmentation possible. However, the dataset does not exhibit multiple different cell types/cell shapes, making it less universal. Cell count: Models can be trained to identify cell number, cell shape, perhaps also identifying dividing cells as well, since there is staining for chromosomes as well as Tubulin/Survivin, which both mark certain stages of division.
3.6 SR-CACO-2 Dataset Limitations
First, it would have been interesting to supplement this data set with oversaturation/undersaturation or very bright vs very dim imaging, as live imaging of cells requires low laser power to minimize damage to cells. This therefore often results in dim images, compared to what can be acquired using fixed and immunostained images. Second, Caco2 cells were ideal for cell for use in this study due to the ease and efficiency of culturing them in 3D and their ability to efficiently polarize. However, often the next step in biological studies is to use human patient-derived tissue or mouse tissue. Future studies could involve the imaging of either primary human or primary mouse organoids. Third, live cells are more dynamic, while fixed cells are static. Live cells are actively growing, dividing, and synthesizing building blocks. This results in changes to the markers that we have used, like mCherry-Histone H2B and GFP-tubulin, which can change shape and position within an individual cell over time, and provide important biological information, such as the stage of cell division. This information cannot be followed for an individual cell once fixed and immunostained, which may impact the SISR model performance. Although, live cells look nearly identical to fixed cells, using fixed dead cells for training SISR models remains a practical solution. Ideally, collecting images of live cells as time-lapse videos for training could be a better choice as it reflects the test time scenario. However, acquiring such data is extremely challenging for the task of SISR due to cell movements or the appearance of new cells due to cell division. This will create a large misalignment between tiles at different scales, rendering the training impractical. Fourth, the dataset is limited in terms of the types of proteins that have been captured. We are only looking at chromosomes, cell membrane/cell structural proteins, and proteins specific for division. The dataset could perhaps be improved with an increase in the number and type of markers used, such as proteins that appear as puncta in the cytoplasm, proteins localized to the apical membrane (mutually exclusive localization with E-cadherin), basement membrane proteins (extracellular matrix found around the outside of cysts), and signalling molecules.
Figure 6: Interpolated vs Microscopic low-resolution image for scale /2 and CELL0. For interpolation, we used bicubic interpolation of the HR image to obtain a low-resolution with a down scale factor of 2.
3.7 Low-Resolution Images: Interpolation vs Microscope
We conducted an empirical analysis to compare both images: real (microscopic) LR vs. bicubic LR. Here are some relevant observations:
  • Pixel-intensity difference: both LR types are different. Their absolute pixel-wise difference value can be in [0, 50] as well as [200, 255] (see the |REAL - BICUBIC| histogram in Fig.6). The difference is mainly concentrated over cell regions.
  • Compared to real LR, bicubic LR can maintain better cell structure and even full cells from the HR images (see the real and bicubic images in Fig.6).
  • Intensity span: depending on the image, the real LR images are sometimes much more expressive in terms of pixel intensity as they span larger intensities compared to bicubic LR (see the pixel-intensity histogram in Fig.6).
  • Real LR images are more noisy compared to bicubic LR ones which tend to be very smooth (see the visual result of the Laplace filter, and the histogram of its absolute value |Laplace LR| in Fig.6). Note that the noise in real LR is the results of a single scan while HR are scanned 9 times to be averaged.
This large contrast between real and bicubic LR only confirms that bicubic LR cannot be used as a replacement to substitute real LR images for training and evaluation, as done in SR methods over natural scene images. Real LR images must be used to effectively simulate a realistic scenario for the model at deployment time. Therefore our collection of real LR images is extremely valuable to designing realistic SR models.
3.8 Experimental Protocol
@@
Patches sampling. Since the imaged tiles are quite large ( ${\leavevmode\nobreak\ 9k\times 9k}$), they are not well suited for machine learning models as they usually operate on smaller images. To this end, we pre-processed the tiles to patches as presented in Fig.4. On HR tiles of the brightest cell, i.e. CELL2, a sliding window was employed that scans from left to right and from top to bottom of the entire tile. Windows can overlap by ${25\%}$ to increase the chance of capturing large patterns and minimize cropping cells. The window size is set to ${576\times 576}$ with 6464{64}64 extra margin at each side that will be useful later for registration. This additional margin was discarded, and only windows of size ${512\times 512}$ were preserved as final patches.
@@
Many of the sampled windows will land on empty regions, i.e., black background. Across all the 22 HR tiles of CELL2, ${74.24\%}$ of total pixels are background while only ${25.74\%}$ are foreground (cells). Windows which are more black will not be helpful in learning or evaluation since they are easy to reconstruct using SISR models. They would introduce an evaluation bias by giving a high level of performance, which does not reflect the true performance of the models to super-resolve relevant regions, i.e., cells. For this reason, we discarded the windows without enough cell content. To detect these windows, we apply a threshold to tile with a value of 44{4}4 which is deemed sufficient to spot the cells without including noise. Only windows with at least ${20\%}$ of cell content were preserved.
@@
Data split. First, the 22 tiles are randomly split into 15 for training, 3 for validation, and 4 for testing. This prevents mixing patches of the same tile across different sets. Tiles with identifiers (ID) = [’9’, ’10’, ’14’, ’20’] are assigned to the test set. Tiles with ID [’7’, ’11’, ’19’] are used for the validation set. Finally, tiles with ID = [’1’, ’2’, ’3’, ’4’, ’5’, ’6’, ’8’, ’12’, ’13’, ’15’, ’16’, ’17’, ’18’, ’21’, ’22’] are used for training. This is summarized in Tab.1 along with the patch distribution for all subsets.
@@ 4 Single-Image Super-Resolution Methods
In addition to the proposed dataset, we provide an extensive benchmarking of different SISR methods. In particular, 16 representative SISR methods were selected from 4 common families [92] are evaluated. We also compare to a simple baseline which is interpolation via the bicubic method. Our results can serve as baselines for future comparisons. We selected the following state-of-the-art methods as follows: Pre-upsampling SR: SRCNN [19], VDSR [42], DRRN [77], and MemNet [78]. Post-upsampling SR: NLSN [59], DFCAN [65], SwinIR [55], EDSR (LIIF) [16], ENLCN [95], GRL [52], ACT [99], and Omni-SR [85]. Iterative up-and-down sampling SR: DBPN [28], and SRFBN [54]. Progressive upsampling SR: ProSR [89], and MS-LapSRN [44].
The post-upsampling family is the most recent and dominant approach [92]. Other families have seen less progress mainly due to their high computational cost.
@@ 4.1 Evaluation Metrics
The evaluation is performed on the full patch, referred to as full image, as commonly done. We furthermore report refined performance over the cells only, referred to as ROI only, to better assess the prediction quality without the black background inside the patches. ROIs are determined by thresholding the HR patch using a set of thresholds [^6] Evaluation thresholds are [4, 5, 6, 7, 8, 9, 10].. A performance is computed per-threshold over the ROI only. The final reported performance is the average of all per-threshold performances. Models are trained on each cell type, and each scale, separately.
@@ 4.2 Implementation Details
SISR task: Tables 4, 5 show more results over all the methods on ROI only, i.e., cells, and on full image, respectively. Figure 8 shows the PSNR performance in a visual form across all the cells, scales, and methods. Note that performance over full images are relatively way higher compared to when only ROI is considered. Overall, SISR methods achieved better results than simple interpolation. As commonly known, the higher the scale factor, the more difficult the task. This is also observed in our results across all methods and performance measures. For instance, SRCNN [19] yields a PSNR of 35.0835.08{35.08}35.08, 33.1733.17{33.17}33.17, and 29.1929.19{29.19}29.19 for X2, X4, and X8, respectively. Another emerging pattern in these results is the discrepancy performance across cell types: CELL0, CELL1, and CELL2. Over PSNR and SSIM performance, CELL0, i.e. dimmest, is the easiest while CELL2 is the most difficult, i.e. the brightest. Over NRMSE, results are mixed. Across all four studied SISR families, we observe that pre-upsampling SR yielded better results overall across all metrics and scales. This is unexpected since post-upsampling SR are state-of-the-art approaches [92]. These results may suggest that upscaling our LR images through deep models may fail to reconstruct the details of our HR images. This may be explained by the nature of the details present in our HR images which paradoxically comes from noise . Confocal microscopes perform multiple scans to produce a set of slices that are aggregated to produce a final 2D image. Due to an inherited noise produced by the detectors on microscopes, each slice is different. However, averaging these slices decreases the noise and maintains the true signal. The work of [83] shows that it is difficult for CNN-based models to reconstruct input images with noise as they tend to filter out that noise even when it is provided as input. Therefore, it is challenging for such models to produce similar images to our HR images. Such results should be considered in designing future SISR methods for this dataset. We provide in Fig.9 typical visual results for scale X2 over CELL2 which illustrate the limited capacity of these methods to produce accurate HR details as they yield blurry images.
In some cases, the methods fail, yielding a decline in performance such as the case of MS-LapSRN [44]. We have analyzed the reason for this decline in performance and found that the model fails to converge properly as the loss does not decrease as expected. This is most clear on CELL0 data leading to a large decline compared to CELL1 and CELL2. We believe that this is caused by the combination of the very low level of brightness for the cell, and the use of residual learning in MS-LapSRN. This model has difficulty producing very small residuals. Instead, it yields large residuals. This can be confirmed when measuring performance over a full image, i.e., by including a dark background where the typical intensity is 0, which leads to a sharp decline in performance. By inspecting some predicted samples, the intensity of these background regions is typically 10. Note that this is not the case for the ProSR model that also uses residual learning. However, ProSR [89] employs a different architecture than MS-LapSRN which, in this case, is much more efficient due to its depth or its local residual layers. During our experiments, we observed that hyper-parameters do not transfer well across cell types or scales of a model. Therefore, we performed a grid search separately for each model, for each cell type, and each scale. This provides each method with a good and fair chance to perform well. Failure cases are mostly related to the method and the data’s nature.
@@
Downstream biology tasks: Object detection/segmentation. To evaluate the image results from each model on downstream biology tasks, we evaluated each model at X2, X4, and X8 using an object segmentation task using widely used and openly accessible software for image analysis (ImageJ) [71] [^7] https://imagej.net/ij. 30 random images from test set are selected. Images were resized to 512 x 512 pixels with no interpolation. Automatic object segmentation was performed using Star-convex Shapes (StarDist) [70, 93, 94] with the following parameters: modelChoice = Versatile (fluorescent nuclei), normalizeInput = TRUE, percentileBottom = 1.1, percentileTop = 100.0, probThresh = 0.6, nmsThresh = 0.4. The resulting output was the number of objects detected. Manual object counting was performed to create the Ground Truth number of objects. Manual comparisons were made to identify oversegmentation and under segmentation errors for each image. Graphing and statistical analysis were performed using JMP Pro (version 17.0.0) [^8] https://www.jmp.com/en_ca/home.html. Steel’s test for non-parametric multiple comparisons using HR as the control group was performed to assess model performance. Our objective was to evaluate model performance relative to HR, not ground truth. Models that perform as well as HR (i.e. not statistically different than HR at p=0.05) are represented in orange. Figures 10, 11 show results for cell detection across the 2 extreme cases: CELL2 and CELL0, respectively. Cell segmentation performance is reported in Fig.12. Overall, several methods (orange) achieved comparable results to HR images on these tasks which is promising. Figure 13 shows an example of cell segmentation of different methods.